OpenStudy (anonymous):

5x-12=13-4x need to understand the steps

OpenStudy (uri):

Same like before..get all x values together and all numericals together.

OpenStudy (uri):

After = sign,sign changes.

OpenStudy (uri):

If you shift them

OpenStudy (uri):

5x-12=13-4x 4x+5x=13+12 Solve this and show.

OpenStudy (anonymous):

Step 1: move all x terms to one side(left side is my choice) and all constant terms to the other side(right side is my choice).

OpenStudy (anonymous):

The general Idea is to move terms around so that one ends up with x = blah blah blah. In this case 9x = 25 How do you isolate x in this case?

OpenStudy (anonymous):

ok what is step 2?? I am writing it down

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

If you divided 9x by 9 what do you end up with?

OpenStudy (anonymous):

x

OpenStudy (uri):

4x+5x=13+12 9x=25 divide 9 with 25 So,25/9

OpenStudy (uri):

x=25/9

OpenStudy (anonymous):

ok I'm to 9x=25

OpenStudy (texaschic101):

now divide by 9 to get x by itself...but remember what you do to one side you have to do to the other side

OpenStudy (anonymous):

Divide both sides by 9. What do you end up with? Can you do this one in your head?

OpenStudy (anonymous):

ok I got it. You are unable to simplify 25/9. lets work on another one

OpenStudy (anonymous):

OK